OpenStudy (anonymous):

write the equation of the line that is parallel to a line containing (2,-5),(7,1) through (5,3) Steps so i can learn or explain

OpenStudy (anonymous):

Alright so this lone you speak of it has to be parallel to another line right

OpenStudy (anonymous):

Line*

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Find the gradient of the original line

OpenStudy (anonymous):

The gradient of the parallel line will be the same

OpenStudy (anonymous):

how to do that ?

OpenStudy (anonymous):

Then you can find the equ. Of line of the parallel line using the coordinates of the point through which the parallel line passes through

OpenStudy (anonymous):

You dont know how to find the gradient of the original line?

OpenStudy (anonymous):

Gradient is sometimes called slope.

OpenStudy (anonymous):

And thus it shall be

OpenStudy (anonymous):

You know how to find slope

OpenStudy (anonymous):

okay i found it to be -4/-5 which will be 4/5 right ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

so it will be y=4/5x+b

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Dude the gradient of the line is 6/5

OpenStudy (anonymous):

Not 4/5

OpenStudy (anonymous):

and 6/5 will be Y=6/5x+b ?

OpenStudy (anonymous):

1-(-5)=6

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

But remember you have to find the value of b too

OpenStudy (anonymous):

how to find the value of b ?

OpenStudy (anonymous):

As i said before pnce you get y=6/5x+b substitute the values of x and y by (5,3)

OpenStudy (anonymous):

Once*

OpenStudy (anonymous):

b should equal 3 ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

my answer should be y=6/5x+ -3b

OpenStudy (anonymous):

Lets hope so

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

Good luck

OpenStudy (anonymous):

Where did the "b" in your equation come from?

OpenStudy (anonymous):

The y intercept

OpenStudy (anonymous):

In other words magic

OpenStudy (anonymous):

b represents the y-intercept when the equation written in slope-intecept form, but it is only a placeholder. The "b" is usually not written in the equation.

OpenStudy (anonymous):

whaaaaa?

OpenStudy (anonymous):

your saying using another variable is used in the place of b

OpenStudy (anonymous):

cause if thats the case i usually use c instead of b

OpenStudy (anonymous):

The answer is \(y=\frac{6}{5}x-3\) not \(y=\frac{6}{5}x-3b\)

OpenStudy (anonymous):

ohh ya actually now that you mention it...

OpenStudy (anonymous):

there should be no b